Fermi National Laboratory


Accelerator Update for October 24, 2005 - October 26, 2005

Monday October 24
The day shift began with Operations monitoring store 4465 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with NuMI off due to chiller work, and with MiniBooNE and the Meson MCenter experiments taking beam.

At 8:49 AM, Operations reported a vacuum burst in the Main Injector (MI) P1 transfer line. Operators recovered the vacuum system.

At 10:19 AM, Operations reported another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system.

Operations began preparing the Recycler at 11:05 AM to stash antiprotons.

The stash was put on hold until Pbar completed a reverse proton tune up at 1:19 PM. Linac technicians then reported that they needed to replace a power supply for VT5IN. The techs completed this job at 1:29 PM.

At 1:24 PM, Operations reported that the first stash of antiprotons was lost in MI. Operators investigated and discovered that a timing event hadn't gone out.

The stash ended and Pbar resumed stacking at 2:11 PM. MCenter resumed normal operations by 2:16 PM.

A MI RF expert reported at 2:20 PM that they were preparing the MIRF stations for a bunch rotation study.

At 4:19 PM, the H- Source vacuum gauge tripped off. Operations notified an expert.

The Pbar bunch rotation study ended at 6:45 PM. Pbar resumed stacking.

NuMI experimenters informed Operations that they had completed their chiller work for the day. Operations established beam to NuMI at 5:05 PM.

At 5:11 PM, Operations reported that a vacuum pump (I:IP7102) tripped off for the first of twenty times throughout the rest of the evening shift. The pump reset each time.

The Recycler began preparing to stash at 6:58 PM. After three transfers, Operations reported that the Recycler had run out of bucket space. Operations contacted an expert who advised them to let the beam cooldown for a bit. Operations did this, the beam shrank, and the last transfers went in.

At 8 PM, Operations reported a Tevatron (TeV) kicker (AAK1) tripped off. Store 4465 remained in the TeV, but CDF and D0 turned off their silicon detectors. An expert came in and charged the kicker. The tripped cleared, but the supply wouldn't turn on. The investigation continued.

The stash ended and Pbar resumed stacking at 8:32 PM. MCenter resumed taking beam.

Operations terminated store 4465 at 10:04 PM.

Operations prepared D0 for a collision hall access at 10:30 PM.

A TeV kicker expert reported that he had completed his repairs on AAK1 at 11:13 PM.

Tuesday October 25
The midnight shift began with Operations monitoring the stash, with TeV in a dry squeeze, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations put the TeV into shot setup at 1:29 PM.

Operations established Recycler only store 4468 at 3:28 AM with an initial luminosity of 144.906E30. This store set a New Luminosity Record.

The Recycler prepared at 3:31 AM to stash.

At 3:53 AM, the P150 line suffered another vacuum burst.

At 4:32 AM, Operations reported that the first Recycler stash fell out of MI. Operators investigated and discovered that the dampers were off.

The Recycler stash ended and Pbar and MCenter resumed normal operations by 5:15 AM.

A NuMI expert reported at 7:32 AM that they were continuing Horn chiller work. No beam to NuMI.

The Recycler began preparing at 9:04 AM to stash. The stash ended and Pbar and MCenter resumed normal operations at 10:15 AM.

A MI expert began an 8-GeV study at 11:41 AM. The study lasted for about 2.5 hours.

A NuMI expert reported at 12:52 PM that the experiment's compressor for their target and Horn was installed and running.

Pbar experts began some AP1 transfer line studies at 1:19 PM.

At 1:36 PM, Operations reported that a number of vacuum pumps tripped off in the P150 transfer line, but there was no vacuum cascade. All the pumps reset.

The Recycler prepared at 2:12 PM to stash.

The Recycler stash ended and Pbar and MCenter resumed normal operations around 3:24 PM.

At 3:34 PM, an interlock problem with the Linac Buncher halted beam. Experts investigated, but could find nothing wrong. The Buncher came back okay, but then LRF2 tripped off on the same interlock indication. The interlock module was reseated and beam resumed about thirty minutes later.

NuMI resumed taking beam at 3:45 PM.

At 4:43 PM, store 4468 aborted when a TeV sector B2 VCB opened. Experts investigated and later replaced a control power supply and a standby power unit.

Cryo system technicians reported at 5:22 PM that they would use this downtime to repair a TeV sector A1 wet engine.

Operations established beam to the Meson MTest experiment at 5:42 PM. They took beam for 1.75 hours and then shutdown for the night.

Operations reported at 6:37 PM that they had started the TeV turn on.

The Recycler began preparing at 9 PM to stash. The Recycler took four transfers, but then had no room for more antiprotons. Pbar and MCenter had resumed normal operations by 10:07 PM.

TeV experts began a study period at 8:31 PM.

Wednesday October 26
The midnight shift began with Operations monitoring the stash, with TeV experts conducting studies, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing at 4:16 AM to stash.

TeV experts completed their studies at 5 AM.

At 5:08 AM, the P150 transfer line suffered a vacuum burst. Operators closed vacuum valves to contain the cascade and then reset everything.

The stash ended. Pbar and MCenter had resumed normal operations by 5:21 AM.

Operators turned off the TeV and then began preparing the TeV, CDF, and D0 at 5:39 AM for access. D0 discovered a preamp power supply in their detector that failed and had to be replaced for them to take data. They must open their detector to do this work. They estimated that this work would take 10 to 12 hours. The TeV and CDF will do maintenance work.

The Plans for Wednesday
The plans for today are to allow TeV maintenance work this morning and then conduct TeV safety system checks this afternoon while D0 repairs its detector. The Booster will conduct some disruptive beam studies. The next shot setup might start later this afternoon or early evening, whenever D0 finishes their work.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/26/2005   email Fermilab